OpenStudy (anonymous):

Solve Each Equation and Check the Solution w/6=-3/4

OpenStudy (johnweldon1993):

w/6 = -3/4 we want w by itself so multiply both sides by 6 what is -3/4 times 6?

OpenStudy (anonymous):

w=-2.57

OpenStudy (zale101):

w=-9/2

OpenStudy (johnweldon1993):

@Zale101 has the right idea

OpenStudy (johnweldon1993):

and check it -4.5/6 = -3/4 -.75 = -.75 indeed it does...so -4.5 is correct